# ZeroTuning: Unlocking the Initial Token’s Power to Enhance Large Language Models Without Training

### v2

@@ Abstract
Token-level attention tuning—a class of training-free methods including Post-hoc Attention Steering (PASTA) and Attention Calibration (ACT)—has emerged as a promising way to improve frozen LLMs with interpretable interventions. However, these methods depend on auxiliary heuristics to identify “important” task-specific tokens, which can introduce bias and limit applicability when token importance is unclear or when using optimized kernels where attention maps are inaccessible. We propose a simpler and more elegant alternative: acting only on the initial token (e.g., <BOS> in LLaMA). We show theoretically that adding lightweight biases to this token’s attention logits monotonically controls the entropy of the downstream attention distribution—an effect amplified by its natural function as an attention sink. Our empirical analysis reveals that this tuning process can positively affect LLMs and better unlock their pretrained knowledge, with stronger effects in early layers and distinct scaling preferences across attention heads. Building on these insights, we introduce ZeroTuning: a training-free method that improves LLM performance by applying head-specific attention adjustments to the initial token, requiring zero parameter updates. We present two variants: a supervised mode that calibrates on validation examples, and a novel unsupervised mode that directly minimizes the model’s output entropy. Our method requires no KV‑cache or decoding changes, and is kernel‑agnostic (works with SDPA and FlashAttention). The method is lightweight and requires only four lines of modification to standard LlamaAttention code. It achieves broad gains across 15 datasets and outperforms previous, more complex methods; for instance, with Llama-3.1-8B, it yields relative improvements of 19.9% on classification, 4.5% on question answering, and 2.1% on dialogue. ZeroTuning also works out-of-the-box with quantized inference and maintains its performance improvements with increasing context lengths. Our code and runnable demo are available at https://anonymous.4open.science/r/ZeroTuning.
@@ 1 Introduction
Training-free methods have been widely explored to enhance Large Language Models (LLMs) performance at inference time. Among these, token-level attention tuning has emerged as a particularly promising direction, offering an interpretable way to steer model behavior by modifying the attention distribution without any parameter updates. Unlike fine-tuning (Hu et al., 2021; Dettmers et al., 2023) or prompt engineering (Wei et al., 2022; Wang et al., 2022; Han et al., 2025a; b), which largely treat LLMs as black boxes, attention tuning provides a transparent mechanism to guide the model’s focus. Methods such as Post-hoc Attention Steering (PASTA (Zhang et al., 2023a), AutoPASTA (Zhang et al., 2024a)), and Attention Calibration (ACT) (Yu et al., 2024) have demonstrated the power of this approach, even outperforming prompting-based techniques in complex tasks like open-domain question answering (Zhang et al., 2024a). Similar principles have been successfully applied to vision-language models to mitigate hallucinations by re-weighting attention towards image tokens (Liu et al., 2024b; Zhu et al., 2024; Wei & Zhang, 2024).
However, the efficacy of these methods is fundamentally constrained by their reliance on external, often heuristic, mechanisms to identify task-specific ”important” tokens. This dependency not only introduces the risk of bias (e.g., amplifying misleading cues) but also limits their applicability in scenarios where token importance is ambiguous or when using optimized attention kernels where attention maps are inaccessible. This critical limitation motivates a fundamental question: Is it possible to enhance model performance by tuning a universal, task-agnostic token, thereby bypassing the need for fragile, task-specific token identification?
@@
In this paper, we demonstrate that the answer is a definitive yes. The solution lies not in adding complexity, but in leveraging a ubiquitous yet often overlooked architectural artifact: the initial token (e.g., <BOS> in LLaMA). While its tendency to function as an ”attention sink” is well-documented (Xiao et al., 2023; Kaul et al., 2024; Gu et al., 2024; Barbero et al., 2025), its potential as an active control lever for performance enhancement has remained largely untapped.
Our theoretical analysis reveals that modulating the attention on this single token allows for precise control – sharpening or smoothing – over the entire downstream attention distribution while preserving the relative importance of task-relevant tokens. Its natural role as an attention sink further amplifies this effect (Sec 3.1). Furthermore, we show this control can be achieved even without direct access to attention weights by modifying the initial token’s key or query states.
To validate this insight, we conducted a systematic investigation and uncovered three key findings:
  1. 1.
    Despite lacking semantic meaning, the initial token acts as a positive and effective control point for steering model behavior. Tuning its attention corrects the model’s biases and unlocks its pretrained knowledge with reduced output entropy. This approach consistently yields greater performance gains than tuning any other token (Sec 3.2).
  2. 2.
    The influence of this tuning is systemic, propagating consistently across layers. The shallow and middle layers are the most impactful, though jointly tuning all layers produces the strongest results (Sec 3.3).
  3. 3.
    The effect is heterogeneous across attention heads; some respond positively to increased attention (up-effective), while others respond negatively (down-effective). We show that selectively targeting the dominant head type outperforms uniform tuning (Sec 3.4,Sec 3.5).
Building on these findings, we introduce ZeroTuning (see Figure 1), a simple, powerful, and training-free method that recalibrates the initial token’s attention to boost LLM performance without any task-specific identification (Sec 3.6). We introduce two variants for attention calibration: a supervised mode by maximizing the accuracy on the labeled validation set, and a novel unsupervised mode by minimizing output entropy. Across a suite of 15 benchmarks, ZeroTuning achieves substantial gains on models like Llama-3.1-8B-Instruct, Llama-2-13B-Instruct, Qwen-2-7B, and Deepseek-R1-14B. For instance, it boosts Llama-3.1-8B-Instruct performance by a relative 19.9% on classification, 4.5% on question answering, and raises its MT-Bench score from 7.804 to 7.966. The method demonstrates remarkable robustness across long contexts, few-shot settings, quantization, and prompt variations. Our work not only delivers a practical tool for lightweight model enhancement but also sheds new light on a fundamental control mechanism within LLMs, advancing both inference-time optimization and model interpretability.
@@ 2 Related Work
Our work is situated at the intersection of two active research areas: inference-time attention tuning and the mechanistic understanding of initial tokens. A growing body of work has shown that modifying token-level attention at inference time can enhance the performance of both LLMs and VLMs (Yu et al., 2024; Zhang et al., 2023a; Liu et al., 2024b; Wei & Zhang, 2024). However, prevailing methods like PASTA (Zhang et al., 2023a) and Auto-PASTA (Zhang et al., 2024a), which identify and up-weight key tokens, or ACT (Yu et al., 2024), which down-weights non-initial sink tokens, fundamentally rely on heuristics to identify task-specific tokens. This reliance limits their universality and introduces potential biases. Concurrently, another line of research has focused on explaining why the initial token often becomes an ”attention sink” (Xiao et al., 2023), attributing it to architectural biases and its role as a stabilizing anchor (Barbero et al., 2025; Gu et al., 2024). While these studies provide a crucial understanding of what the phenomenon is, the question of how to actively and elegantly harness it for performance gains remains largely unexplored. Our work bridges this gap. We shift the focus from task-specific token identification to a universal, task-agnostic control point, and move from passive observation of the initial token to a practical tuning framework that leverages its unique properties. Detailed related work is provided in Appendix A.
@@ 3 Unveiling the Power of the Initial Token
In this section, we first formalize the mechanism of tuning the initial token’s attention, then empirically demonstrate its unique importance and dissect its effects across the model’s layers and heads. This systematic analysis culminates in our proposed ZeroTuning methodology. Unless otherwise specified, all experiments use the Llama-3.1-8B-Instruct model, with setup details in Section 4.1.
@@ 3.1 Formalizing the Tuning Process
In a decoder-only Transformer, autoregressive generation for a sequence $\mathbf{X}=[x_{0},x_{1},\dots,x_{T-1}]\in\mathbb{R}^{d\times T}$ involves a causal self-attention mechanism. At timestep $T$, the query is derived from the final token representation, $x_{T-1}$, which attends to all preceding token representations (including itself) as keys. This process yields an attention weight distribution over the input sequence:
@@
Intuitively, $\gamma>1$ amplifies $a_{0}$, flattening the remaining distribution, while $\gamma<1$ suppresses $a_{0}$, sharpening it.
Theoretically, the magnitude of this effect is governed by the initial token’s own attention weight, $a_{0}$. We define this effect, $E_{\text{diff},i,j}$, as the change in attention difference between any two non-initial tokens $i$ and $j$:
@@
To analyze how $E_{\text{diff},i,j}$ varies with $a_{0}$, we take its partial derivative with respect to $a_{0}$:
@@
Given that $|a_{i}-a_{j}||\gamma-1|\geq 0$ and the denominator $((\gamma-1)a_{0}+1)^{2}=D^{2}>0$, the derivative is non-negative. Thus, $E_{\text{diff},i,j}$ is a monotonically non-decreasing function of $a_{0}$. In any non-trivial case ($\gamma\neq 1$ and $a_{i}\neq a_{j}$), the relationship is strictly increasing. A detailed proof and visualization are in Appendix C.
This result provides a crucial insight: the larger the initial token’s attention score ($a_{0}$), the more powerful it becomes as a lever for controlling the entire attention distribution. Since prior work has established that initial tokens are natural attention sinks (Barbero et al., 2025), they are inherently potent control points for this tuning process.
@@ 3.2 The Unique Importance of the Initial Token
Given the special role of the initial token, we first investigate a key empirical question: (a) Does tuning its attention positively impact performance on downstream tasks? (b) Is this position more effective and influential than others? To investigate, we conduct a controlled experiment in which we uniformly scale the attention scores of a single token position across all heads and layers using a scaling factor $\gamma$. We evaluate the resulting performance on three downstream tasks: SST-2, BoolQ, and LogiQA. For comparison, we repeat the same procedure for other positions, including the second, third, middle ($\lfloor T/2\rfloor$), and final tokens. As shown in Figure 2, tuning the attention of the initial token consistently yields the largest and most stable performance gains across all tasks. Interestingly, performance varies with the direction of tuning: tasks like SST-2 benefit from up-scaling ($\gamma>1$), while BoolQ and LogiQA improve with down-scaling ($\gamma<1$).
Figure 3: The impact of the tuning.
Previous work has identified the initial token as an attention sink that helps prevent over-mixing of information during autoregressive generation (Gu et al., 2024; Barbero et al., 2025). Our empirical results extend this understanding, showing that tuning the initial token’s attention beneficially reshapes the distribution over subsequent tokens and, in turn, improves model performance across tasks. This beneficial effect can be understood from two perspectives.
1. Correcting Biases. The tuning process counteracts reasoning flaws caused by pretrained attention biases for each dataset, as visualized in Figure 3. Our error analysis reveals a clear pattern: (a) Up-scaling ($\gamma>1$) benefits tasks requiring holistic context integration. It flattens the attention distribution, preventing the model from over-focusing on misleading local cues. For instance, in SST-2 sentiment analysis (see Appendix F), LLMs often fixate on isolated negative keywords while ignoring the surrounding positive context. Increasing the initial token’s attention promotes a more balanced, global understanding, thereby correcting such biased predictions. (b) Down-scaling ($\gamma<1$) benefits tasks demanding sharp focus on critical evidence. It amplifies the relative weights of non-initial tokens, helping the model pinpoint key details in a diffuse context. In long-context tasks like BoolQ, a model’s focus can become diluted, failing to locate the precise text segment containing the answer. Reducing the initial token’s attention sharpens the model’s focus on this salient information.
@@ 3.3 Layer-wise Analysis of Initial Token Scaling
To understand the propagation of this effect, we examine how its effect varies when applied selectively across different layers. Following prior work on layer functionality in transformer-based models (Jin et al., 2024; Zhang et al., 2024b), we divide the 32 layers of Llama-3.1-8B-Instruct into three groups: shallow (Layers 1–10), middle (Layers 11–21), and deep (Layers 22–31). We then perform independent attention scaling experiments for each group on six tasks: BoolQ, SST-2, SST-5, MR, LogiQA, and MathQA. Based on earlier findings, we apply a scaling range of $[0,1]$ for BoolQ and LogiQA, and $[1,2]$ for the remaining tasks.
@@
As shown in Figure 5, tuning yields consistent trends at any depth (i.e., accuracy changes similarly with scaling regardless of depth), and jointly tuning all layers amplifies these benefits, often resulting in the highest accuracy. However, the magnitude of improvement varies. In most cases, tuning the shallow and middle layers leads to greater accuracy than tuning the deep layers.
@@ 3.4 Analyzing the Role of the Initial Token Across Attention Heads
Unlike layers passing information sequentially, attention heads operate in parallel and contribute independently via concatenation. It remains unclear how they differ in response to the initial token. To investigate this, we increase the initial token’s attention of each head individually by applying $\gamma=1.5$, and evaluate the model’s performance on SST-2 and MMLU. For comparison, we also evaluate (i) no scaling ($\gamma=1$), (ii) uniform up-scaling ($\gamma=1.5$) across all heads, and (iii) uniform down-scaling ($\gamma=0.6$) across all heads.
As shown in Figure 6, attention heads exhibit distinct behaviors in response to initial token amplification. We categorize heads as up-effective if this modification improves performance, and down-effective if it results in performance degradation. Interestingly, the relative proportions of up-effective and down-effective heads vary across datasets, which in turn explains the observed differences in response to uniform scaling. For example, SST-2 contains more up-effective heads and thus benefits from uniform up-scaling. In contrast, MMLU has a higher proportion of down-effective heads, making down-scaling more effective than up-scaling.
@@ 3.6 ZeroTuning Methodology
Building on our empirical findings, we propose ZeroTuning, a method that enhances LLM performance via head-specific attention adjustments to the initial token, without requiring task-specific token identification. The core methodology involves three steps: head behavior profiling, selective rescaling, and renormalization. To accomplish the first two steps, hyperparameter calibration, we introduce both supervised and unsupervised variants to cater to different application scenarios.
@@
To mitigate the reliance on labeled data, we propose a novel unsupervised calibration strategy. Based on our finding that a model’s output entropy strongly correlates with its accuracy (as detailed in Section 3.2), this variant bypasses the calibration set entirely. Instead, it identifies the optimal heads and scaling factor $\gamma$ by minimizing the average next-token prediction entropy directly on the unlabeled test inputs.
Finally, the third step, Renormalization, is applied in both variants by re-normalizing the scaled attention scores via the softmax function to maintain a valid probability distribution. For optimized attention implementations (e.g., SDPA, Flash Attention) where direct score modification is infeasible, ZeroTuning applies scaling to the query or key states. We demonstrate and show that this yields similar effects in Appendix G.

### v3

@@ Abstract
Token-level attention tuning—a class of training-free methods including Post-hoc Attention Steering (PASTA) and Attention Calibration (ACT)—has emerged as a promising approach for improving frozen LLMs via interpretable interventions. However, these methods rely on auxiliary heuristics to identify “important” task-specific tokens, which can introduce bias and limit applicability when token importance is ambiguous or when optimized kernels make attention maps inaccessible. We propose a simpler alternative: intervening only on the initial token (e.g., <BOS> in LLaMA). We theoretically show that adding lightweight biases to this token’s attention logits systematically shifts and reshapes downstream attention patterns—an effect amplified by its natural role as an attention sink. Empirically, we find that this tuning can improve LLM performance and better elicit pretrained knowledge, with stronger effects in early layers and distinct scaling preferences across attention heads. Building on these findings, we introduce ZeroTuning, a training-free method that improves LLM performance by applying head-specific attention adjustments to the initial token, requiring no parameter updates. We present two variants: a supervised mode that calibrates on validation examples, and an unsupervised mode that directly minimizes output entropy. ZeroTuning requires no KV-cache or decoding changes and is kernel-agnostic (works with SDPA and FlashAttention). It requires only four lines of modification to standard LlamaAttention code, achieves gains across 15 datasets, and outperforms prior, more complex methods. For example, on Llama-3.1-8B it yields relative improvements of 19.9% on classification, 4.5% on question answering, and 2.1% on dialogue. ZeroTuning also works out of the box with quantized inference and maintains its improvements as context length increases. Our work provides a lightweight tool for inference-time improvement, advancing both optimization and interpretability. Our code and runnable demo are available at https://anonymous.4open.science/r/ZeroTuning.
@@ 1 Introduction
Training-free methods have been widely explored to enhance the performance of Large Language Models (LLMs) at inference time. Among these, token-level attention tuning has emerged as a particularly promising direction, offering an interpretable way to steer model behavior by modifying attention distributions without any parameter updates. Unlike fine-tuning (Hu et al., 2021; Dettmers et al., 2023) or prompt engineering (Wei et al., 2022; Wang et al., 2022; han2025readthinkmitigatingllm), which largely treat LLMs as black boxes, attention tuning provides a transparent mechanism for guiding the model’s focus. Methods such as Post-hoc Attention Steering (PASTA (Zhang et al., 2023a), AutoPASTA (Zhang et al., 2024a)) and Attention Calibration (ACT) (Yu et al., 2024) have demonstrated the potential of this approach, in some cases outperforming prompting-based techniques on complex tasks such as open-domain question answering (Zhang et al., 2024a). Similar principles have also been applied to vision-language models to mitigate hallucinations by re-weighting attention toward image tokens (Liu et al., 2024b; Zhu et al., 2024; Wei and Zhang, 2024).
However, the efficacy of these methods is fundamentally constrained by their reliance on external, often heuristic, mechanisms for identifying task-specific “important” tokens. This dependency not only introduces the risk of bias (e.g., amplifying misleading cues) but also limits applicability when token importance is ambiguous or when optimized attention kernels make attention maps inaccessible. This limitation motivates a fundamental question: Is it possible to enhance model performance by tuning a universal, task-agnostic token, thereby bypassing the need for fragile, task-specific token identification?
In this paper, we show that the answer is yes. The key is not added complexity, but rather leveraging a ubiquitous yet often overlooked architectural artifact: the initial token (e.g., <BOS> in LLaMA). While its tendency to function as an “attention sink” is well documented (Xiao et al., 2023; Kaul et al., 2024; Gu et al., 2024; Barbero et al., 2025), its potential for performance enhancement has remained largely untapped.
@@
Empirically, we find that steering only the initial token’s attention can consistently improve LLM performance across diverse tasks, often matching or surpassing prior attention-steering approaches that rely on tuning multiple, task-specific tokens.
To understand why this simple intervention works, we proceed in three steps: we first provide a theoretical characterization of how scaling the initial token reshapes attention (Sec. 3.1), then connect the induced redistribution to changes in systematic error patterns and output entropy (Sec. 3.2), and finally localize where the control is most effective by dissecting its impact across layers and heads (Sec. 3.33.5)
Insights. Steering the BOS token provides a universal, training-free method to improve LLMs: it modulates attention sharpness, reduces output entropy, mitigates understanding biases, and concentrates the effect in early layers and specific heads.
Building on these findings, we introduce ZeroTuning (Figure 1), a simple and training-free method that recalibrates the initial token’s attention to boost LLM performance without task-specific token identification (Sec. 3.6). We present two variants: a supervised mode that maximizes accuracy on a labeled validation set, and an unsupervised mode that minimizes output entropy. Across 15 benchmarks, ZeroTuning yields substantial gains on models including Llama-3.1-8B-Instruct, Llama-2-13B-Instruct, Qwen-2-7B, and Deepseek-R1-14B. For example, it improves Llama-3.1-8B-Instruct by a relative 19.9% on classification and 4.5% on question answering, and increases its MT-Bench score from 7.804 to 7.966. The method is robust across long contexts, few-shot settings, quantization, and prompt variations.
@@ 2 Related Work
Our work is situated at the intersection of two active research areas: inference-time attention tuning and the mechanistic understanding of initial tokens. A growing body of work has shown that modifying token-level attention at inference time can enhance the performance of both LLMs and VLMs (Yu et al., 2024; Zhang et al., 2023a; Liu et al., 2024b; Wei and Zhang, 2024). However, prevailing methods like PASTA (Zhang et al., 2023a) and Auto-PASTA (Zhang et al., 2024a), which identify and up-weight key tokens, or ACT (Yu et al., 2024), which down-weights non-initial sink tokens, fundamentally rely on heuristics to identify task-specific tokens. This reliance limits their universality and introduces potential biases. Concurrently, another line of research has focused on explaining why the initial token often becomes an ”attention sink” (Xiao et al., 2023), attributing it to architectural biases and its role as a stabilizing anchor (Barbero et al., 2025; Gu et al., 2024). While these studies provide a crucial understanding of what the phenomenon is, the question of how to actively and elegantly harness it for performance gains remains largely unexplored. Our work bridges this gap. We shift the focus from task-specific token identification to a universal, task-agnostic control point, and move from passive observation of the initial token to a practical tuning framework that leverages its unique properties. Detailed related work is provided in Appendix A.
@@ 3 Unveiling the Power of the Initial Token
In this section, we formalize how scaling the initial token’s attention reshapes the attention distribution, then empirically demonstrate why this token is uniquely effective and analyze how the effect varies across layers and heads. These findings motivate our ZeroTuning methodology. Unless otherwise specified, all experiments use Llama-3.1-8B-Instruct; see Section 4.1 for setup details.
@@ 3.1 Formalizing the Tuning Process
In a decoder-only Transformer, autoregressive generation for a sequence $\mathbf{X}=[x_{0},x_{1},\dots,x_{T-1}]\in\mathbb{R}^{d\times T}$ relies on causal self-attention. At timestep $T$, the query derived from the final token representation $x_{T-1}$ attends to all preceding token representations (including itself) as keys, producing an attention distribution:
@@
Intuitively, $\gamma>1$ amplifies $a_{0}$, flattening the remaining distribution, while $\gamma<1$ suppresses $a_{0}$, sharpening it. The magnitude of this effect is governed by the initial token’s own attention weight, $a_{0}$. We define $E_{\text{diff},i,j}$ as the change in the attention difference between any two non-initial tokens $i$ and $j$:
@@
To analyze how $E_{\text{diff},i,j}$ varies with $a_{0}$, we take the partial derivative with respect to $a_{0}$:
@@
Since $|a_{i}-a_{j}||\gamma-1|\geq 0$ and the denominator $((\gamma-1)a_{0}+1)^{2}=D^{2}>0$, the derivative is non-negative. Thus, $E_{\text{diff},i,j}$ is monotonically non-decreasing in $a_{0}$; in the non-trivial case ($\gamma\neq 1$ and $a_{i}\neq a_{j}$), it is strictly increasing. A detailed proof and visualization are in Appendix C.
Key insight: the larger the initial token’s attention ($a_{0}$), the more effectively it can act as a lever to control the attention distribution. Since initial tokens are known to be attention sinks (Barbero et al., 2025), they are natural control points for this tuning process.
@@ 3.2 The Unique Importance of the Initial Token
Given the special role of the initial token, we first investigate a key empirical question: (a) Does tuning its attention positively impact performance on downstream tasks? (b) Is this position more effective and influential than others? To investigate, we conduct a controlled experiment in which we uniformly scale the attention scores of a single token position across all heads and layers using a scaling factor $\gamma$. We evaluate the resulting performance on three downstream tasks: SST-2, BoolQ, and LogiQA. For comparison, we repeat the same procedure for other positions, including the second, third, middle ($\lfloor T/2\rfloor$), and final tokens. As shown in Figure 2, tuning the initial token consistently yields the largest and most stable gains across tasks. Interestingly, the best direction depends on the task: SST-2 benefits from up-scaling ($\gamma>1$), while BoolQ and LogiQA improve with down-scaling ($\gamma<1$).
Previous work has identified the initial token as an attention sink that helps prevent over-mixing of information during autoregressive generation (Gu et al., 2024; Barbero et al., 2025). Our results extend this understanding by showing that tuning the initial token reshapes attention over subsequent tokens and improves downstream performance. We interpret this effect from two perspectives.
@@
1. Correcting Biases. Tuning counteracts reasoning failures driven by pretrained attention biases. We observe a consistent pattern: (a) Up-scaling ($\gamma>1$) benefits tasks requiring holistic context integration. It flattens attention over non-initial tokens, reducing over-reliance on misleading local cues. For example, in SST-2 (Appendix F), models can fixate on isolated negative words while missing positive context; increasing initial-token attention promotes a more global read and corrects such errors. (b) Down-scaling ($\gamma<1$) benefits tasks requiring sharp focus on critical evidence. It increases the relative contrast among non-initial tokens, helping the model localize salient spans in diffuse contexts. In long-context BoolQ, reducing initial-token attention can sharpen focus on the text segment that contains the answer.
@@ 3.3 Layer-wise Analysis of Initial Token Scaling
To understand how this effect propagates, we apply scaling selectively to different layer groups. Following prior work on layer functionality in Transformers (Jin et al., 2024; Zhang et al., 2024b), we partition the 32 layers of Llama-3.1-8B-Instruct into shallow (Layers 1–10), middle (Layers 11–21), and deep (Layers 22–31). We then scale the initial token within each group and evaluate on six tasks: BoolQ, SST-2, SST-5, MR, LogiQA, and MathQA. Based on Section 3.2, we use $[0,1]$ for BoolQ and LogiQA, and $[1,2]$ for the remaining tasks.
@@
As shown in Figure 4, scaling induces similar trends across depths, and tuning all layers jointly typically yields the best performance. However, the gains vary in magnitude: tuning shallow and middle layers usually helps more than tuning deep layers.
@@ 3.4 Analyzing the Role of the Initial Token Across Attention Heads
Unlike layers, attention heads operate in parallel and contribute independently via concatenation, so their responses to initial-token scaling may differ. To test this, we scale the initial token in each head individually with $\gamma=1.5$ and evaluate on SST-2 and MMLU. For comparison, we also consider (i) no scaling ($\gamma=1$), (ii) uniform up-scaling across all heads ($\gamma=1.5$), and (iii) uniform down-scaling across all heads ($\gamma=0.6$).
As shown in Figure 5, heads respond heterogeneously to initial-token amplification. We call a head up-effective if scaling improves performance, and down-effective otherwise. The mix of these head types varies across datasets, which helps explain why uniform scaling behaves differently: SST-2 has more up-effective heads and benefits from up-scaling, whereas MMLU has more down-effective heads and thus prefers down-scaling.
@@ 3.6 ZeroTuning Methodology
Building on these findings, we propose ZeroTuning, which improves LLM performance by applying head-specific scaling to the initial token without identifying task-specific tokens. ZeroTuning has three steps: (i) head behavior profiling, (ii) selective rescaling, and (iii) renormalization. For calibration (steps (i)–(ii)), we introduce both supervised and unsupervised variants.
@@
To mitigate the reliance on labeled data, we propose a novel unsupervised calibration strategy. Based on our finding that a model’s output entropy strongly correlates with its accuracy (Section 3.2), this variant does not require any labeled calibration or validation set. Instead, it identifies the optimal heads and scaling factor $\gamma$ by minimizing the average next-token prediction entropy over a batch of unlabeled inputs. Crucially, the unlabeled inputs can be obtained in two practical ways: (i) in an offline setting, by running ZeroTuning on a held-out, unlabeled corpus from the same domain; or (ii) in a test-time adaptation style, by performing entropy-based search over the current batch of test-time queries. We provide a more detailed analysis of this unsupervised variant in Appendix E.
Finally, Renormalization is shared by both variants: we re-normalize the scaled scores (via softmax) to maintain a valid distribution. For optimized attention implementations (e.g., SDPA, FlashAttention) where direct score modification is infeasible, ZeroTuning applies scaling to the query or key states; Appendix G shows this yields similar effects.